find the following values for the function defined by f(x)=x+4

x x+4 f(x) (x,y)

2 ------ ----- -------

type in order of pair in the third input box

Answer by jim_thompson5910(35256)

Notice how if x=2, then x+4=2+4=6. So x+4=6 when x=2. This then means that f(2)=6




x                x+4       f(x)        (x,y)


2                 6        6           (2, 6)




If you need more help, email me at jim_thompson5910@hotmail.com

Also, feel free to check out my tutoring website

Jim